Calculate the first AND second derivative dy/dx and d^2y/dx^2 for the curve given by:

x(t)=t-t^2, y(t)=3t-t^2
Answer #1

a(t) = t-t² = 1-2t at y (t) = 3t-t² dy 3- 2t c at dy 3-22 3-2t delat dalat 2 dy dx dx 1-2t 1-2t and 3-24 3-26 at d24 dx2 el dplease rate after understanding the solution..have a good day